# A comprehensive survey on deep active learning and its applications in medical image analysis

### v1

@@ 4.4 Region-based Active Learning: Smaller Labeling Unit
Most AL works require the oracle to label the full image in medical image analysis and computer vision. However, labeling a full image can introduce redundancy in fine-grained tasks like segmentation and detection, resulting in an inefficient use of the annotation budget. For example, in segmentation tasks of autonomous driving, large areas in the image (e.g., roads) do not need exhaustive annotation. Instead, those annotation budgets would be better spent on detailed smaller areas, such as pedestrians or utility poles. An image can be divided into non-overlap regions to further improve annotation efficiency, and experts can opt to annotate specific regions within an image. This method is termed as “region-based active learning”. This section introduces region-based active learning from three perspectives: patches, superpixels, and region-based active domain adaptation.
@@ 4.4.1 Patches
Patches are most commonly used in region-based active learning, generally represented as square boxes. Mackowiak et al. (2018) combined uncertainty and annotation cost to select the informative patches for annotation. Casanova et al. (2020) employed deep reinforcement learning to automatically select informative patches for annotation. In retinal blood vessels segmentation of eye images, Xu et al. (2021) selected patches with the highest uncertainty for annotation. Furthermore, they utilized latent-space Mixup to encourage linearization between labeled and unlabeled samples, thus leveraging unlabeled data to improve performance.
@@ 4.4.2 Superpixels
Superpixels are also widely used in region-based active learning. Superpixel generation algorithms (Achanta et al., 2012; Van den Bergh et al., 2012) over-segment images based on color and texture, grouping similar pixels into the same superpixel. Superpixel-based AL initially pre-segments the images and then calculates the informativeness of each superpixel. The informativeness metric of each superpixel is the average of its constituent pixels. Kasarla et al. (2019) introduced a baseline method for selecting superpixels based on uncertainty. In multi-view indoor scene segmentation, Siddiqui et al. (2020) adopted uncertainty and disagreement between different viewpoints to select informative superpixels for annotation. Cai et al. (2021) utilized uncertainty as the informativeness metric. They introduced a class-balanced sampling strategy to better select superpixels containing minority classes. Furthermore, they adopted a “dominant labeling” scheme. The dominant labeling is the majority class label of all pixels in the superpixel. They assigned the dominant labeling to every pixel within a superpixel, thus eliminating the need for detailed delineation. Results showed that, with the same number of labeling clicks, dominant labeling at the superpixel level significantly outperforms precise labeling at the patch level. As a follow-up, Kim et al. (2023) proposed to adaptively merge and split spatially adjacent, similar, and complex superpixels, respectively. This approach yielded better performance than (Cai et al., 2021) with dominant labeling. In 3D vision, similar over-segmentation has also been applied to point clouds segmentation by Wu et al. (2021a), Hu et al. (2022) and Liu et al. (2022a). The former two works adopted supervoxel segmentation algorithms, while the latter employed k-Means for over-segmentation.
4.4.3 Region-based Active Domain Adaptation
To better utilize the annotation budget, some ADA segmentation works also employed patches or superpixels for annotation. In Xie et al. (2022a), uncertainty and regional impurity were used to select and annotate the most informative patches. Regional impurity measured the number of unique predicted classes within the neighborhood of a pixel, which presents the edge information. They used extremely small patches (e.g., size of 3x3) for annotation and achieved performance close to full supervision with only 5% of the annotation cost. Wu et al. (2022b) proposed a density-based method to select the most representative superpixels in the target domain for annotation. They employed Gaussian mixture models (GMM) as density estimators for superpixels in both the source and target domains, aiming to select those with high density in the target domain and low density in the source domain.
@@ 4.5.1 Synthetic Samples as Data Augmentation
The simplest approach considers the synthetic sample produced by generative models as advanced data augmentation. These methods utilize label-conditioned generative models. As a result, it’s guaranteed that all synthetic samples are correctly labeled since specifying the labels is a prerequisite for data generation. This method enables us to acquire more labeled samples without any additional annotations. Tran et al. (2019) argued that most synthetic samples produced by generative models are not highly informative. Therefore, they first adopted the BALD uncertainty to select samples for annotation, then trained a VAE-ACGAN on these labeled data to generate more informative synthetic samples. Mahapatra et al. (2018) used conditional GANs to generate chest X-rays with varying diseases to augment the labeled dataset. Then, MC Dropout was used to select and annotate highly uncertain samples. With the help of AL and synthetic samples, they achieved performance near fully supervised using only 35% of the data. Training conditional generative models require a large amount of labeled data, while the labeled dataset in AL is often relatively small. To address this issue, Lou et al. (2023) proposed a conditional SinGAN (Shaham et al., 2019) that only requires one pair of images and masks for training. The SinGAN improved the annotation efficiency for nuclei segmentation. Additionally, Chen et al. (2022b) integrated implicit semantic data augmentation (ISDA) (Wang et al., 2021) into AL. They initially used ISDA to augment unlabeled samples, then selected samples with large diversity between different data augmentations for annotation. The model is trained on both the original data and its augmentations.
@@ 5 Active Learning for Medical Image Analysis
Due to the potential of significantly reducing annotation costs, AL is receiving increasing attention in medical image analysis. The unique traits of medical imaging require us to design specialized AL methods. Building on the foundation of the previous two sections, this section will focus on introducing AL works tailored to medical image analysis across different tasks, including classification, segmentation, and reconstruction. Additionally, in Table 3, we list all the AL works related to medical image analysis in this survey, providing the name of the used dataset, its modality, ROIs, and corresponding clinical and technical tasks.
Year Venues Modality ROIs Dataset Clinical Task Technical Task
Zhou et al. (2017) 2017 CVPR Colonoscopy Colon in-house Image Quality Assessment Classification
Colonoscopy Colon in-house Polyp Detection Classification
CT Lung in-house Pulmonary Embolism Detection Classification
Gal et al. (2017) 2017 ICML Dermscopy Skin ISIC 2016 Skin Cancer Diagnosis Classification
Yang et al. (2017) 2017 MICCAI Histopathology Colon Glas Gland Segmentation Segmentation
Ultrasound Lymoh Node in-house Lymoh Node Segmentation Segmentation
Beluch et al. (2018) 2018 CVPR Fundus Eye EyePacs Diabetic Retinopathy Detection Classification
Xu et al. (2018) 2018 CVPR Histopathology Colon Glas Gland Segmentation Segmentation
Sourati et al. (2018) 2018 DLMIA MRI Brain dHCP & in-house Brain Extraction Segmentation
Sourati et al. (2019) 2019 TMI
Kuo et al. (2018) 2018 MICCAI CT Head in-house Intracranial Hemorrhage Detection Segmentation
Mahapatra et al. (2018) 2018 MICCAI X-ray Chest SCR & Chestx-ray8 Lung Segmentation Segmentation
@@
Jin et al. (2019) 2019 arXiv MRI Heart Cardiac Atlas Project MRI Reconstruction Reconstruction
Knee fastMRI MRI Reconstruction Reconstruction
@@
Zheng et al. (2019) 2019 AAAI Histopathology Colon Glas Gland Segmentation Segmentation
@@
MRI Heart HVSMR 2016 Whole-heart Segmentation Segmentation
Electron Microscopy Fungus in-house Fungus Segmentation Segmentation
Zhang et al. (2019) 2019 CVPR MRI Knee fastMRI MRI Reconstruction Reconstruction
@@
Qi et al. (2019) 2019 JBHI Histopathology Breast BreaKHis Breast Cancer Diagnosis Classification
@@
Sadafi et al. (2019) 2019 MICCAI Mircoscopy Blood in-house Red Blood Cell Detection Object Detection
Zheng et al. (2020) 2020 AAAI MRI Heart HVSMR 2016 Whole-heart Segmentation Segmentation
Electron Microscopy Mouse Lee et al. (2015) Neuron Boundary Segmentation Segmentation
Lin et al. (2020) 2020 ECCV Electron Microscopy Mouse Synapses & Mitochondria EM-R50 Synapse Detection & Mitochondria Segmentation Segmentation
Dai et al. (2020) 2020 MICCAI MRI Brain BraTS 2019 Brain Tumor Segmentation Segmentation
Li and Yin (2020) 2020 MICCAI Histopathology Colon Glas Gland Segmentation Segmentation
MRI Brain iSeg Infant Brain Segmentation Segmentation
Liu et al. (2020) 2020 MICCA CT Lung DeepLesion Pulmonary Nodule Detection Object Detection
Mi et al. (2020) 2020 MICCAI Electron Microscopy Mouse Cortex SNEMI3D Accelerated Acquisition of Electron Microscopy Reconstruction
Human Cerebrum in-house Reconstruction
Pineda et al. (2020) 2020 MICCAI MRI Knee fastMRI MRI Reconstruction Reconstruction
Shen et al. (2020) 2020 MICCAI Immunohistochemistry Breast in-house Breast Cancer Region Segmentation Segmentation
Wang et al. (2020a) 2020 MICCAI CT Lung Tianchi Lung Diease Detection Classification
Fundus Eye EyePacs Diabetic Retinopathy Detection Classification
Bakker et al. (2020) 2020 NeurIPS MRI Knee fastMRI MRI Reconstruction Reconstruction
Brain fastMRI MRI Reconstruction Reconstruction
Hiasa et al. (2020) 2020 TMI CT Hip & Thigh TCIA & in-house Muscle Segmentation Segmentation
Table 3: Surveyed Works of Active Learning related to Medical Image Analysis.
Year Venues Modality ROIs Dataset Clinical Task Technical Task
Huang et al. (2020) 2020 TMI X-ray Chest in-house Rib Fracture Recognition Object Detection
Zhao et al. (2021) 2021 JBHI Dermscopy Skin ISIC 2017 Skin Lesion Segmentation Segmentation
X-ray Hand RSNA Bone Age Dataset Finger Bone Segmentation Segmentation
Wu et al. (2021b) 2021 MedIA CT Lung CC-CCII COVID-19 Diagnosis Classification
Zhou et al. (2021b) 2021 MedIA Colonoscopy Colon in-house Image Quality Assessment Classification
Colonoscopy Colon in-house Polyp Detection Classification
CT Lung in-house Pulmonary Embolism Detection Classification
Wang and Yin (2021) 2021 MICCAI Mircoscopy (Synthetic) Bacterial Cells VGG Cell Cell Counting Keypoint Localization
Histopathology Human Bone Marrow MBM Cell Counting Keypoint Localization
Histopathology Human Adipocyte Cells ADI Cell Counting Keypoint Localization
- Various Tissues & Species DCC Cell Counting Keypoint Localization
Xu et al. (2021) 2021 MICCAI Fundus Eye DRIVE Retina Vessel Segmentation Segmentation
OCTA Eye ROSE-1 Retina Vessel Segmentation Segmentation
Zhou et al. (2021a) 2021 MICCAI CT Lung MSD Tumor Segmentation Segmentation
Colon MSD Tumor Segmentation Segmentation
Kidney KiTS 19 Kidney & Tumor Segmentation Segmentation
Nguyen et al. (2021) 2021 MIDL X-ray Chest in-house Diagnosis of Airspace Opacity & Lung Lesion Classification
Chest RSNA Pneumonia Diagnosis of Pneumonia Classification
Chest CheXpert Detection of Pleural Effusion Classification
Mahapatra et al. (2021) 2021 TMI X-ray Chest ChestX-ray8 Thoracic Disease Diagnosis Classification
Histopathology Colon Glas Gland Segmentation Segmentation
Nath et al. (2021) 2021 TMI CT Pancreas MSD Pancreas & Tumor Segmentation Segmentation
MRI Hippocampus MSD Hippocampus Segmentation Segmentation
Chen et al. (2021) 2021 TPAMI CT Heart in-house Calcification Level Prediction in Aortic Stenosis Classification
Wang et al. (2022a) 2022 arXiv CT Lung AAPM CT Reconstruction Reconstruction
Spine VerSe CT Reconstruction Reconstruction
Kothawade et al. (2022b) 2022 AAAI X-ray Chest PneumoniaMNIST Pneumonia & Normal Classification Classification
Quan et al. (2022) 2022 CVPR X-ray Head Kaggle Cephalometric Landmark Detection Keypoint Localization
Hand Payer et al. (2019) Hand Landmark Detection Keypoint Localization
Zhang et al. (2022a) 2022 CVPR MRI Spine in-house Diagnosis of Metastatic Epidural Spinal Cord Compression Classification
@@
Jin et al. (2022c) 2022 Knowledge-based Systems Dermscopy Skin ISIC 2020 Skin Lesion Classification Classification
Jin et al. (2022b) 2022 Knowledge-based Systems Dermscopy Skin ISIC 2018 Skin Lesion Segmentation Segmentation
X-ray Chest Jaeger et al. (2013) Lung Segmentation Segmentation
Atzeni et al. (2022) 2022 MedIA MRI Brain SATA Brain Structure Segmentation Segmentation
Histology Brain in-house Brain Structure Segmentation Segmentation
Dai et al. (2022) 2022 MedIA MRI Brain BraTS 2019 Brain Tumor Segmentation Segmentation
Brain MALC Brain Structure Segmentation Segmentation
Zhou et al. (2022) 2022 MedIA CT Lung MSD Tumor Segmentation Segmentation
CT Colon MSD Tumor Segmentation Segmentation
CT Kidney KiTS 19 Kidney & Tumor Segmentation Segmentation
Colonoscopy Colon CVC-ClinicDB Polyp Segmentation Segmentation
Year Venues Modality ROIs Dataset Clinical Task Technical Task
Nath et al. (2022) 2022 MICCAI CT Liver MSD Liver & Tumor Segmentation Segmentation
Hepatic Vessels MSD Hepatic Vessels & Tumor Segmentation Segmentation
Bai et al. (2022) 2022 MICCAI Wireless Capsule Endoscopy Colon CAD-CAP Polyp Segmentation Segmentation
Balaram et al. (2022) 2022 MICCAI X-ray Chest Chestx-ray8 Thoracic Disease Diagnosis Classification
Wu et al. (2022c) 2022 MICCAI CT Liver LiTS Liver & Tumor Segmentation Segmentation
CT, MRI Liver CHAOS Liver & Tumor Segmentation Segmentation
CT Liver Sliver07 Liver & Tumor Segmentation Segmentation
CT Liver MSD Liver & Tumor Segmentation Segmentation
Kothawade et al. (2022c) 2022 MICCAIW X-ray Chest PneumoniaMNIST Pneumonia & Normal Classification Classification
Histopathology Colon PathMNIST Survival Prediction Classification
Mircoscopy Peripheral Blood BloodMNIST Cell Type Classification Classification
Dermscopy Skin ISIC 2018 Skin Lesion Diagnosis Classification
@@
Fundus Eye APTOS-2019 Diabetic Retinopathy Detection Classification
Li et al. (2022) 2022 TMI Histopathology Prostate PANDA Gleason Grading of Prostate Cancer Classification
Mahapatra et al. (2022) 2022 TMI X-ray Chest ChestXpert Thoracic Disease Diagnosis Classification

### v3

@@ 4.3 Region-based Active Learning: Smaller Labeling Unit
Most AL works require the oracle to label the full image in medical image analysis. However, labeling a full image can introduce redundancy in fine-grained tasks like segmentation or detection, resulting in an inefficient use of the annotation budget. For the example of abdomen multi-organ segmentation, large organs that are easy to segment (e.g., liver or spleen) do not need exhaustive annotation. Instead, those budgets would be better spent on small organs that are hard to segment, like the esophagus and adrenal glands. To address this issue, images could be divided into non-overlap regions for higher annotation efficiency, and experts can opt to annotate specific regions within an image, which is termed “region-based active learning”. This section introduces region-based active learning from the perspectives of patches and superpixels, which means that AL methods mentioned in this section selected either patches or superpixels within an image for annotation.
@@ 4.3.1 Patches
Patches are most commonly used in region-based active learning, generally represented as square boxes. Mackowiak et al. (2018) combined uncertainty and annotation cost to select the informative patches for annotation. In retinal blood vessels segmentation of fundus images, Xu et al. (2021) selected patches with the highest uncertainty for annotation. Furthermore, they utilized latent-space mixup to encourage linearization between labeled and unlabeled samples, thus leveraging unlabeled data to improve performance. Casanova et al. (2020) employed deep reinforcement learning to automatically select informative patches for annotation. In grey matter and white matter segmentation of pathology images, Lai et al. (2021) first split the whole slide image into multiple patches. With the confidence (i.e., maximum predictive probability) of each patch, a mean filter of size 5x5 is used to aggregate the confidence of the neighboring patches. As a result, one aggregated metric corresponded to a region of 5x5 patches, and regions with the highest uncertainty were selected for annotation. Besides, Qiu et al. (2023) adopted an adaptative region selection with non-square patches for whole-slide images. Instead of sampling square patches, they dynamically determined the size of each non-square patch by carefully locating an informative area on each slide. The proposed method demonstrated improvement in annotation efficiency and robustness to AL hyperparameters compared to the square patch baseline.
@@ 4.3.2 Superpixels
Superpixels are also widely used in region-based active learning. Superpixel-based AL initially pre-segments the images with superpixel generation algorithms based on color and texture (Achanta et al., 2012; Van den Bergh et al., 2012), and then calculates the informativeness of each superpixel. The informativeness metric of each superpixel is the average of its constituent pixels. Siddiqui et al. (2020) adopted uncertainty and disagreement between different viewpoints to select informative superpixels for annotation. In OCT segmentation, Kadir et al. (2023) proposed edge-based entropy and divergence to select highly uncertain superpixels for annotation. Experiments on three datasets were conducted to illustrate the effectiveness of their method. For superpixel selection, Cai et al. (2021) proposed dominant labeling which is the majority class label of all pixels in the superpixel. They assigned the dominant labeling to every pixel within a superpixel, thus eliminating the need for detailed delineation. They further introduced a class-balanced sampling strategy to better select superpixels containing minority classes. Results showed that dominant labeling with superpixels significantly outperforms precise labeling with patches under the same number of labeling clicks. As a follow-up work, Kim et al. (2023) proposed to adaptively merge and split spatially adjacent, similar, and complex superpixels, respectively. This approach yielded better performance than Cai et al. (2021) with dominant labeling. Li et al. (2023) utilized the superpixels to estimate the regional consistency which is the difference between the prediction and the dominant class of each superpixel. Combining other metrics like entropy and diversity, they selected the most uncertain foreground and background superpixel to reduce the annotation cost.
@@ 4.4.1 Synthetic Samples as Data Augmentation
The simplest approach considers the synthetic sample produced by generative models as advanced data augmentation. These methods utilize label-conditioned generative models. As a result, it’s guaranteed that all synthetic samples are correctly labeled since specifying the labels is a prerequisite for data generation. This method enables us to acquire more labeled samples without any additional annotations. Tran et al. (2019) argued that most synthetic samples produced by generative models are not highly informative. Therefore, they first adopted the BALD uncertainty to select samples for annotation, then trained a VAE-ACGAN on these labeled data to generate more informative synthetic samples. Mahapatra et al. (2018) used conditional GANs to generate chest X-rays with varying diseases to augment the labeled dataset. Then, MC Dropout was used to select and annotate highly uncertain samples. With the help of AL and synthetic samples, they achieved performance near fully supervised using only 35% of the data. Training conditional generative models requires a large amount of labeled data, while the labeled dataset in AL is often relatively small. To address this issue, Lou et al. (2023) proposed a conditional SinGAN (Shaham et al., 2019) that only requires one pair of images and masks for training. The SinGAN improved the annotation efficiency for nuclei segmentation. Chen et al. (2022b) integrated implicit semantic data augmentation (ISDA) (Wang et al., 2021) into AL. They initially used ISDA to augment unlabeled samples, then selected samples with large diversity between different data augmentations for annotation. The model is trained on both the original data and its augmentations. Mahapatra et al. (2024) trained a VAE for synthesizing informative and non-redundant samples. These samples are generated by first sampling in the latent space of VAE and feeding them to the VAE decoder. Besides, scores of label preservation and redundancy avoidance were adopted to pick the most informative synthetic samples. The proposed method was tested in chest X-ray classification and multiple toy datasets from MedMNIST (Yang et al., 2023).
@@ 4.4.2 Generative Active Learning
4.5 Active Domain Adaptation: Tackling Distribution Shift
Domain Adaptation (DA) (Guan and Liu, 2021) has wide applications in medical image analysis. It aims to transfer knowledge from the source to the target domain, thus minimizing annotation costs. Currently, the most common setting of DA is unsupervised domain adaptation (UDA), in which the source domain is labeled while the target domain is unlabeled. For the example of abdominal multi-organ segmentation, we can train a domain-adaptive segmentation model with labeled MR images alongside unlabeled CT images to achieve good performance on the CT domain (Liu et al., 2023b). However, the performance of UDA still lags behind fully supervised learning in the target domain. Selecting and annotating informative samples would be beneficial to bridge this gap. This setting is known as active domain adaptation (ADA). For better queries in ADA, one should consider both uncertainty and representativeness regarding the target domain. The latter is commonly referred to as domainness or targetness in ADA. This section reviews the image-wise and region-wise ADA.
4.5.1 Image-wise Active Domain Adaptation
In this section, ADA methods performed an image-level selection, which involves most of the ADA works. Su et al. (2020) was the first to introduce the concept of ADA and combined domain adversarial learning with AL. Through a domain discriminator and task model, they performed importance sampling to select target domain samples that are uncertain and highly different from the source domain. Fu et al. (2021) combined query-by-committee, uncertainty, and domainness in ADA. They adopted a domain discriminator to select samples with high domainness and employed Gaussian kernels to filter out anomalous and source-similar samples of the target domain. Random sampling was also used to improve diversity. Prabhu et al. (2021) performed k-Means clustering on target domain samples and selected cluster centers for annotation. The cluster centers were weighted by uncertainty, thus ensuring that selected samples were uncertain and diverse. In segmentation tasks, Ning et al. (2021) introduced the idea of anchors in ADA. They concatenated features of different classes from the source domain images. Cluster centers of these concatenations were referred to as anchors. They then computed the distance between each target sample and its nearest anchor. Target samples with the highest distance were requested for annotation. Xie et al. (2022b) introduced the concept of energy (LeCun et al., 2006) into ADA. The energy is inversely proportional to the likelihood of the data distribution. In this work, the model trained on the source domain was used to calculate the energy of target domain samples. Samples with high energy were selected for annotation, which suggested they are representative of the target domain and substantially different from the source data. Huang et al. (2023) selected samples with high uncertainty and prediction inconsistency to their nearest prototypes. In the context of medical image analysis, Chen et al. (2023a) tackled domain shifts in the setting of federated active learning. They proposed an EDL-based framework with a global model across all clients and local models for each client. In this work, the EU is related to domain shifts between the global model and local data. Therefore, the AUs of the global and local models were calibrated by the EU, thus improving performance. Results on multiple medical imaging datasets showed its effectiveness in reducing annotation costs. In nasopharyngeal carcinoma tumor segmentation, Wang et al. (2023) proposed a source-domain and target-domain dual-reference strategy to select informative samples for annotation. Specifically, the features of the source samples were clustered and the cluster centers were reference samples. Target samples with the highest and lowest similarity to the references are selected for annotation, which were treated as domain-invariant and domain-specific samples, respectively.
4.5.2 Region-wise Active Domain Adaptation
To better utilize the annotation budget, some ADA works also selected patches or superpixels within an image for annotation. Shin et al. (2021) proposed LabOR, which first used a UDA pre-trained model to generate pseudo-labels for target samples, which was used to train two segmentation heads. They maximized the disagreements between the two heads and annotated regions that exhibited the most disagreement. LabOR achieved performance close to full supervision with only 2.2% of target domain annotations. In Xie et al. (2022a), uncertainty and regional impurity were used to select and annotate the most informative patches. Regional impurity measured the number of unique predicted classes within the neighborhood of a pixel, which presents the edge information. They used extremely small patches (e.g., size of 3x3) for annotation and achieved performance close to full supervision with only 5% of the annotation cost. Wu et al. (2022b) proposed a density-based method to select the most representative superpixels in the target domain for annotation. They employed Gaussian mixture models (GMM) as density estimators for superpixels in both the source and target domains, aiming to select those with high density in the target domain and low density in the source domain.
@@ 5 Active Learning for Medical Image Analysis
Due to the potential of significantly reducing annotation costs, AL is receiving increasing attention in medical image analysis. The unique traits of medical imaging require us to design specialized AL methods. Building on the foundation of the previous two sections, this section will focus on introducing AL works tailored to medical image analysis across different tasks, including classification, segmentation, and reconstruction.
Additionally, in Table 3, we list all the AL works related to medical image analysis in this survey, providing the name of the used dataset, its modality, ROIs, and corresponding clinical and technical tasks.
Year Venues Modality ROIs Dataset Sampling Unit Size (Train/Val/Test) or (Train+Val/Test) Initial Pool Size Budget per Round Clinical Task Technical Task
Gorriz et al. (2017) 2017 arXiv Dermscopy Skin ISIC 2017 image 2,000 (1,600/400) 600 35 Skin Cancer Diagnosis Classification
Zhou et al. (2017) 2017 CVPR Colonoscopy Colon in-house image 4,000 (2,000/2,000) - - Image Quality Assessment Classification
Colonoscopy Colon in-house image 28,250 (16,300/11,950) Polyp Detection Classification
CT Lung in-house PE candidate 6,255 (3,840/2,415) Pulmonary Embolism Detection Classification
Gal et al. (2017) 2017 ICML Dermscopy Skin ISIC 2016 image 400 (200/200) 100 20 Skin Cancer Diagnosis Classification
Yang et al. (2017) 2017 MICCAI Histopathology Colon GlaS patch 165 images (80/5/80) 8 8 Gland Segmentation Segmentation
Ultrasound Lymoh Node in-house image 74 (37/37) Lymoh Node Segmentation Segmentation
Otálora et al. (2017) 2017 MICCAIW Fundus Eye e-ophtha patch 20,148 (17,520/656/1,972) 160 32 Exudate Classification Classification
Beluch et al. (2018) 2018 CVPR Fundus Eye EyePacs image 88,702 (67,961/3,000/17,741) 1,000 5,000 Diabetic Retinopathy Detection Classification
Xu et al. (2018) 2018 CVPR Histopathology Colon GlaS patch 165 images (80/5/80) 8 8 Gland Segmentation Segmentation
Sourati et al. (2018) 2018 DLMIA MRI Brain dHCP patch 66 patients 3 patinets 50 Brain Extraction Segmentation
Sourati et al. (2019) 2019 TMI in-house 25 patients
Kuo et al. (2018) 2018 MICCAI CT Head in-house region, image, volume 1,247 volumes (934/313) 1/32 double Intracranial Hemorrhage Detection Segmentation
Mahapatra et al. (2018) 2018 MICCAI X-ray Chest SCR & Chestx-ray8 image 647 (247/400) 10% 5% Lung Segmentation Segmentation
@@
Zheng et al. (2019) 2019 AAAI Histopathology Colon GlaS patch 1,530 0 30% & 50% Gland Segmentation Segmentation
Electron Microscopy Fungus in-house 784 Fungus Segmentation Segmentation
@@
MRI Heart HVSMR 2016 slice 20 volumes (10/10) every 2, 10, 20,40, 80 slices Whole-heart Segmentation Segmentation
@@
Qi et al. (2019) 2019 JBHI Histopathology Breast BreaKHis image See paper for details Breast Cancer Diagnosis Classification
@@
Sadafi et al. (2019) 2019 MICCAI Mircoscopy Blood in-house image 208 (188/20) 30 - Red Blood Cell Detection Object Detection
@@
Shi et al. (2019) 2019 MICCAWI Dermscopy Skin ISIC 2017 image 4,332 (3,582/150/600) 10% 10% Skin Lesion Diagnosis Classification
Gu et al. (2018) 2019 TBME Endomicroscopy Breast Gu et al. (2017) image 1,366 (1,042/324) - - Endomicroscopy Mosaic Classification Classification
Colonoscopy Colon Ye et al. (2016) 7,847 (3,947/3,900) - - Gastrointestinal Image Classification
Zheng et al. (2020) 2020 AAAI MRI Heart HVSMR 2016 slice 20 volumes (10/10) 0 every 5, 10, 20, 40, 80 slices Whole-heart Segmentation Segmentation
Electron Microscopy Mouse Lee et al. (2015) slice 4 volumes (3/1) 0 every 4, 16, 64 slices Neuron Boundary Segmentation Segmentation
Lin et al. (2020) 2020 ECCV Electron Microscopy Mouse Synapses EM-R50 image 48.7K (28.7K/20K) - 1,280 Synapse Detection Segmentation
Mouse Synapses & Mitochondria 15K (10K/5K) Mitochondria Segmentation Segmentation
Dai et al. (2020) 2020 MICCAI MRI Brain BraTS 2019 volume 335 patients (260/75) 10 10 Brain Tumor Segmentation Segmentation
slice 500 500
Li and Yin (2020) 2020 MICCAI Histopathology Colon GlaS 2D patch 27,200 10% 20% Gland Segmentation Segmentation
MRI Brain iSeg 3D patch 16,380 - 50% Infant Brain Segmentation Segmentation
Liu et al. (2020) 2020 MICCA CT Lung DeepLesion volume 1,281 (1,000/281) 10% 10% Pulmonary Nodule Detection Object Detection
Shen et al. (2020) 2020 MICCAI Immunohistochemistry Breast in-house image 3,441 (2,767/306+368) 10% 10% Breast Cancer Region Segmentation Segmentation
Wang et al. (2020b) 2020 MICCAI CT Lung Tianchi image 7K (3.5K/3.5K) 5% about 7% Lung Diease Detection Classification
@@
Fundus Eye EyePacs image 4,460 (2,230/2,230) 10% about 11% Diabetic Retinopathy Detection Classification
Hiasa et al. (2020) 2020 TMI CT Hip & Thigh TCIA & in-house slice 20 volumes 5% 5% Muscle Segmentation Segmentation
Huang et al. (2020) 2020 TMI X-ray Chest in-house region 10,966 images (7:1:2) 0% 5% & 10% Rib Fracture Recognition Object Detection
Shen et al. (2021) 2021 ISBI MRI Brain BraTS2018 slice 285 volumes (233:52) 5 volumes 100 & 200 Brain Tumor Segmentation Segmentation